For the diprotic weak acid H2A, Kal =2.0 x 10- and Ko e =8.6 x 10 What is the pH of a 0.0750 M solution of H,A? pH What are t please give an explanation/ show work! thank you!
Answer #1


To find the pH of diprotic weak acid Given that Ku = 2.0x106 K = 8.6x109 0.0750 M solution of H, A It will dissociates HA HA

To find the [HA] [H,A]=0.0750- x [H,A]=0.0750 - 0.000387 = 0.0746M For the secondreaction HAA? +H* Ka = 8.6x109 Equilibrium e